OpenStudy (anonymous):

can you explain what the Van der Waals and ideal gas law is? in the easiest way? thanks!:)

OpenStudy (anonymous):

An ideal gas in one in which molecules have a property such that they no longer interact with each other and all that the kinetic energy will contribute the total energy of the ideal gas. So at pressures bellow than 40 standard atmospheres, there could be shown that the equation of state for such a gas has the simple form PV=NKT, where N stands for the total number of particles in the gas. Do you need any more?

OpenStudy (anonymous):

do you know of a situation that would require you to use the van der waals equation of states?

OpenStudy (anonymous):

See this: http://en.wikipedia.org/wiki/Van_der_Waals_equation

OpenStudy (anonymous):

thanks!

OpenStudy (ybarrap):

Van Der Waals accounts for the fact that atoms take up space and can also attract each other. The ideal gas law assumes neither. In the link that @Saeeddiscover provided $$ \left(p + \frac{n^2 a}{V^2}\right)\left(V-nb\right) = nRT $$ The term after \(p\), the pressure term adjusts for the atom's mutual attraction or repulsion, while the second term after \(V\), the volume term, reduces the volume by a factor \(nb\) and so tries to more accurately model gasses.